# DeepSSM: an emulator of gravitational wave spectra from sound waves during cosmological first-order phase transitions

### v1

@@ 1 Introduction
Recent observations from pulsar timing arrays in the nano-hertz range [1, 2, 3, 4, 5] have provided preliminary evidence for the existence of a stochastic gravitational wave background (SGWB) and have offered initial insights into the shape of its power spectrum. The next generation of space- and ground-based gravitational wave observatories [6, 7, 8, 9, 10, 11] are expected to thoroughly characterise this diffuse signal across different frequency bands. The significant interest in this new observable arises from its connection not only to astrophysical phenomena but also to various sophisticated physical processes in the early Universe. Many of these processes are closely related to physics beyond the Standard Model (BSM) of particle physics [12, 13, 14], highlighting the critical importance of precise SGWB measurements for exploring potential new physics.
Cosmological first-order phase transitions (FOPTs) are some of the most important physical processes that occur during the radiation-dominated era. During these transitions, GWs can be generated through three primary mechanisms: bubble collisions, sound waves, and turbulence. These mechanisms collectively produce diffused GWs, which constitute a significant component of the SGWB. Among these production channels, recent studies [15, 16, 14] suggest that sound waves generally serve as the dominant source of GWs in the case of thermal FOPTs. Models quantifying the resulting GW spectra from the FOPTs are typically based on some phase transition parameters, such as the strength, the duration, the characteristic temperature of the transition, and the bubble wall velocity. Notably, these parameters also depend on the particular extension of the Standard Model, as shown by [17, 18, 19, 20, 21, 22, 23, 24, 25]. Therefore, phase transition parameters serve as crucial links between new physics and the SGWB, which is a physical observable.
@@
Although these semi-analytical models offer a convenient framework for predicting SGWB spectra, they typically require high-precision, multi-dimensional numerical integrations. This results in significant computational overhead, making them impractical to be applied to some analysis pipelines of the SGWB data. In practice, studies such as [37, 38, 39] often resort to empirical templates, such as broken or double-broken power-law models, to perform Bayesian inference, even though there are discrepancies between the results of the semi-analytical model and these power-law models [40]. The situation becomes even more challenging when considering the enhanced SSM, as it introduces additional phase transition parameters and higher-dimensional integrations, further adding computational complexity while producing spectra deviating more significantly from a power-law shape.
Given the necessity of a highly efficient model for computing the GW spectra, developing an emulator for gravitational waves sourced by phase transitions becomes crucial. The goal of this emulator is to efficiently map a set of phase transition parameters to the corresponding GW spectrum. Since this is a highly non-linear projection, neural networks (NNs) present a particularly promising solution. NNs are well-established for their ability to approximate any smooth function [41, 42, 43, 44], making them particularly well-suited to build such emulators. In recent years, NN-based emulators have been developed for computing various cosmological signals, including those from the cosmic microwave background [45, 46, 47, 48, 49, 50] (including temperature, polarisation, and lensing), large-scale structures [51, 52, 53, 54, 55, 56, 57], and 21-cm signals [58, 59, 60]. A similar framework could thus be adopted to predict the SGWB spectra arising from cosmological first-order phase transitions.
@@ 2 GW production from sound waves
where we have defined ~$\tilde{\ell}_{ij}(\tau,\mathbf{k})\equiv h_{ij}(\tau,\mathbf{k})/a(\tau)$ and $\tilde{\Pi}_{ij}(\tau,\mathbf{k})$ is the transverse-traceless (TT) part of the anisotropic stress tensor with $G$ being the gravitational constant. By defining the TT projection operator $\Lambda_{ij,lm}\equiv P_{il}P_{jm}-\frac{1}{2}P_{ij}P_{lm}$, with $P_{ij}\equiv\delta_{ij}-\hat{k}_{i}\hat{k}_{j}$, the term $\bar{\rho}\Pi_{ij}(\tau,\mathbf{k})$ can be expressed as
@@
~$\tilde{\ell}_{ij}(\tau,\mathbf{k})=6\mathcal{H}_{*}\begin{cases}{\displaystyle\int_{\tau_{*}}^{\tau}\frac{d\tau_{1}}{\tau_{1}}}\tilde{\Pi}_{ij}(\tau_{1},\mathbf{k}){\displaystyle\frac{\sin k(\tau-\tau_{1})}{k}},&\tau_{*}<\tau<\tau_{\rm fin},\\ &\\ {\displaystyle\int_{\tau_{*}}^{\tau_{\rm fin}}\frac{d\tau_{1}}{\tau_{1}}}\tilde{\Pi}_{ij}(\tau_{1},\mathbf{k}){\displaystyle\frac{\sin k(\tau-\tau_{1})}{k}},&\tau>\tau_{\rm fin}.\end{cases}$ (2.4)
@@
Therefore, to derive the GW spectrum, one needs to calculate the two-point correlation function of ~$\tilde{\ell}_{ij}^{\prime}$. In the comoving momentum space, for $\tau_{1,2}\gg\tau_{\rm fin}$, we have
~$\begin{split}\langle\tilde{\ell}_{ij}^{~{}^{\prime}}(\tau_{1},\mathbf{k})\tilde{\ell}_{ij}^{*^{\prime}}(\tau_{2},\mathbf{k}_{2})\rangle=&(6\mathcal{H}_{*})^{2}\int_{\tau_{*}}^{\tau_{\rm fin}}\frac{d\tau_{1}}{\tau_{1}}\int_{\tau_{*}}^{\tau_{\rm fin}}\frac{d\tau_{2}}{\tau_{2}}\cos k(\tau-\tau_{1})\cos k_{2}(\tau-\tau_{2})\\ &\times\Big{\langle}\tilde{\Pi}_{ij}(\tau_{1},\mathbf{k})\tilde{\Pi}_{ij}(\tau_{2},\mathbf{k}_{2})\Big{\rangle},\end{split}$ (2.7)
@@
According to eq. (2.10), the key to obtaining the GW spectrum lies in the shear stress, $\tilde{\Pi}_{ij}$, which is determined by the the energy-momentum tensor $T_{ij}$. In the case of a FOPT, the fluid and the scalar field contribute to $T_{ij}$. Therefore, we have
@@ 2.1 The sound shell model
[…]
@@
    Here $\beta$ is the inverse duration of phase transition, $\tilde{T}\equiv\beta T_{i}^{n}$ is the scaled bubble lifetime, and $R_{*}\equiv(8\pi)^{1/3}v_{w}/\beta$ denotes the mean bubble separation, where $v_{w}$ is the bubble wall velocity obtained by evaluating the friction acting on the wall [66, 67, 68]. The bubble lifetime, $\nu(\tilde{T})$, is derived based the above assumptions for the destruction of the bubble and some key quantities describing the dynamics of phase transition (e.g. the bubble nucleation rate, bubble wall velocity, etc.), see details in ref. [30]. In this work, we only consider the exponential nucleation criterion, which means $\nu(\tilde{T})=\exp(-\tilde{T)}$.
@@
[…]
@@ 3 The NN architecture and the training strategy
The architecture of the NN used in our emulator is shown in Fig. 1. It comprises multiple layers of fully connected neurons. The input layer accepts $4$ phase transition parameters $v_{w}$, $\alpha$, $\mathcal{H}_{*}\tau_{\rm sw}$, $\mathcal{H}_{*}R_{*}$. The output layer generates the complete spectrum of $\tilde{\Omega}_{\mathrm{GW}}(K)$, sampled at $100$ logarithmic evenly spaced points in the dimensionless $K$–space, ranging from $1\times 10^{-3}$ to $3\times 10^{2}$. Each of the two hidden layers contains 512 neurons. The projection corresponding to this NN is strictly linear until an activation function is introduced. We adopt the Rectified Linear Unit (ReLU) as the activation function. Both the NN architecture and the activation function are selected based on our experimental results, which demonstrate their superior accuracy compared to other commonly adopted alternatives. It is worth noting that the NN does not take the phase transition temperature $T_{*}$ or the number of degrees of freedom $g_{*}$ as inputs, as the resulting gravitational wave spectrum today depends linearly on these parameters, as predicted by the enhanced SSM.
@@ 4.1 Estimating LISA noise and generating mock data
×$\times\left(1+\left(\frac{f}{8\rm\,mHz}\right)^{4}\right)\left(\frac{1}{2\pi f}\right)^{4}\left(\frac{2\pi f}{c}\right)^{2}.$ (4.3)
They dominate at higher and lower frequencies, respectively, with $P\sim 15$, $A\sim 3$ as noise parameters. Based on this model, the total noise in the measured energy fraction of GWs can be expressed as:
@@
The simulated signals $\Omega_{\rm GW}(f)$, used as test cases, are generated based on the enhanced SSM based on two representative sets of benchmark parameters. The benchmarks, $\mathrm{BP}_{1}$ and $\mathrm{BP}_{2}$, are listed in Table 2. They represent relatively strong phase transitions that occur during the electroweak epoch, resulting in a power-law-like signal and a comparatively flatter signal, respectively (see the blue curves in Fig. 3). The differences in their spectral shapes are mainly attributed to their distinct bubble wall velocities $v_{w}$. We generate the injected mock data using the GW spectra corresponding to these two benchmark points, calculated with the enhanced SSM code rather than the emulator. Any foreground signals, such as those introduced by galactic binaries, are approximately ignored in this work, as they are typically weaker compared to the GW spectra produced by our test cases and can be subtracted through yearly modulation [70]. However, more detailed studies are required when considering some weaker GW spectra introduced by the FOPTs.

### v2

@@ 1 Introduction
Recent observations from pulsar timing arrays in the nano-hertz range [1, 2, 3, 4, 5] have provided preliminary evidence for the existence of a stochastic gravitational wave background (SGWB) and have offered initial insights into the shape of its power spectrum. The next generation of space- and ground-based gravitational wave observatories [6, 7, 8, 9, 10, 11] are expected to thoroughly characterise this stochastic signal across different frequency bands. The significant interest in this new observable arises from its connection not only to astrophysical phenomena but also to various complex physical processes in the early Universe. Many of these processes are closely related to physics beyond the Standard Model (BSM) of particle physics [12, 13, 14], highlighting the critical importance of precise SGWB measurements for exploring potential new physics.
Cosmological first-order phase transitions (FOPTs) are some of the most important physical processes that occur during the radiation-dominated era. During these transitions, GWs can be generated through three primary mechanisms: bubble collisions, sound waves, and turbulence. These mechanisms collectively produce stochastic GWs, which constitute a significant component of the SGWB. Among these production channels, recent studies [15, 16, 14] suggest that sound waves generally serve as the dominant source of GWs in the case of thermal FOPTs. Models quantifying the resulting GW spectra from the FOPTs are typically based on some phase transition parameters, such as the strength, the duration, the characteristic temperature of the transition, and the bubble wall velocity. Notably, these parameters also depend on the particular extension of the Standard Model, as shown by [17, 18, 19, 20, 21, 22, 23, 24, 25]. Therefore, phase transition parameters serve as crucial links between new physics and the SGWB, which is a physical observable.
@@
Although these semi-analytical models offer a convenient framework for predicting SGWB spectra, they typically require high-precision, multi-dimensional numerical integrations. This results in significant computational overhead, making them impractical to be applied to some analysis pipelines of the SGWB data. In practice, studies such as [37, 38, 39, 40] often resort to empirical templates, such as broken or double-broken power-law models, to perform Bayesian inference, even though there are discrepancies between the results of the semi-analytical model and these power-law models [41]. The situation becomes even more challenging when considering the enhanced SSM, as it introduces additional phase transition parameters and higher-dimensional integrations, further adding computational complexity while producing spectra deviating more significantly from a power-law shape.
Given the necessity of a highly efficient model for computing the GW spectra, developing an emulator for gravitational waves sourced by phase transitions becomes crucial. The goal of this emulator is to efficiently map a set of phase transition parameters to the corresponding GW spectrum. Since this is a highly non-linear projection, neural networks (NNs) present a particularly promising solution. NNs are well-established for their ability to approximate any smooth function [42, 43, 44, 45], making them particularly well-suited to build such emulators. In recent years, NN-based emulators have been developed for computing various cosmological signals, including those from the cosmic microwave background [46, 47, 48, 49, 50, 51, 52] (including temperature, polarisation, and lensing), large-scale structures [53, 54, 55, 56, 57, 58, 59, 60], and 21-cm signals [61, 62, 63]. A similar framework could thus be adopted to predict the SGWB spectra arising from cosmological first-order phase transitions.
@@ 2 GW production from sound waves
where we have defined $\tilde{\ell}_{ij}(\tau,\mathbf{k})\equiv h_{ij}(\tau,\mathbf{k})/a(\tau)$ and $\tilde{\Pi}_{ij}(\tau,\mathbf{k})$ is the transverse-traceless (TT) part of the anisotropic stress tensor with $G$ being the gravitational constant. By defining the TT projection operator $\Lambda_{ij,lm}\equiv P_{il}P_{jm}-\frac{1}{2}P_{ij}P_{lm}$, with $P_{ij}\equiv\delta_{ij}-\hat{k}_{i}\hat{k}_{j}$, the term $\bar{\rho}\Pi_{ij}(\tau,\mathbf{k})$ can be expressed as
@@
$\tilde{\ell}_{ij}(\tau,\mathbf{k})=6\mathcal{H}_{*}\begin{cases}{\displaystyle\int_{\tau_{*}}^{\tau}\frac{d\tau_{1}}{\tau_{1}}}\tilde{\Pi}_{ij}(\tau_{1},\mathbf{k}){\displaystyle\frac{\sin k(\tau-\tau_{1})}{k}},&\tau_{*}<\tau<\tau_{\rm fin},\\ &\\ {\displaystyle\int_{\tau_{*}}^{\tau_{\rm fin}}\frac{d\tau_{1}}{\tau_{1}}}\tilde{\Pi}_{ij}(\tau_{1},\mathbf{k}){\displaystyle\frac{\sin k(\tau-\tau_{1})}{k}},&\tau>\tau_{\rm fin}.\end{cases}$ (2.4)
@@
Therefore, to derive the GW spectrum, one needs to calculate the two-point correlation function of $\tilde{\ell}_{ij}^{\prime}$. In the comoving momentum space, for $\tau_{1,2}\gg\tau_{\rm fin}$, we have
$\begin{split}\langle\tilde{\ell}_{ij}^{~{}^{\prime}}(\tau_{1},\mathbf{k})\tilde{\ell}_{ij}^{*^{\prime}}(\tau_{2},\mathbf{k}_{2})\rangle=&(6\mathcal{H}_{*})^{2}\int_{\tau_{*}}^{\tau_{\rm fin}}\frac{d\tau_{1}}{\tau_{1}}\int_{\tau_{*}}^{\tau_{\rm fin}}\frac{d\tau_{2}}{\tau_{2}}\cos k(\tau-\tau_{1})\cos k_{2}(\tau-\tau_{2})\\ &\times\Big{\langle}\tilde{\Pi}_{ij}(\tau_{1},\mathbf{k})\tilde{\Pi}_{ij}(\tau_{2},\mathbf{k}_{2})\Big{\rangle},\end{split}$ (2.7)
@@
According to eq. (2.10), the key to obtaining the GW spectrum lies in the shear stress, $\tilde{\Pi}_{ij}$, which is determined by the energy-momentum tensor $T_{ij}$. In the case of a FOPT, the fluid and the scalar field contribute to $T_{ij}$. Therefore, we have
@@ 2.1 The sound shell model
    $f(\chi)\equiv\frac{4\pi}{\chi}\int_{0}^{\infty}d\xi v_{\rm ip}(\xi)\sin(\chi\xi),\quad l(\chi)\equiv\frac{4\pi}{\chi}\int_{0}^{\infty}d\xi\lambda_{\rm ip}(\xi)\sin(\chi\xi)\,$, (2.21)
@@
    $P_{v}(k)=\frac{k^{2}}{2\pi^{2}\beta^{6}R_{*}^{3}}\int_{0}^{\infty}d\tilde{T}\nu(\tilde{T})\tilde{T}^{6}|\mathcal{A}(\tilde{T}k/\beta)|^{2}.$ (2.23)
@@
$\tilde{\Omega}_{\rm GW}=K^{3}(\bar{P}_{v}^{\rm max})^{2}\int_{0}^{\infty}dPP^{2}\zeta_{v}(P)\int_{-1}^{1}(1-z^{2})^{2}\frac{\zeta_{v}(\tilde{P})}{\tilde{P}^{4}}\Delta(\tau_{\rm sw},R_{*},K,P,\tilde{P})dz,$ (2.25)
@@ 3 The NN architecture and the training strategy
The architecture of the NN used in our emulator is shown in Fig. 1. It comprises multiple layers of fully connected neurons. The input layer accepts $4$ phase transition parameters $v_{w}$, $\alpha$, $\mathcal{H}_{*}\tau_{\rm sw}$, $\mathcal{H}_{*}R_{*}$. The output layer generates the complete spectrum of $\tilde{\Omega}_{\mathrm{GW}}(K)$, sampled at $100$ logarithmic evenly spaced points in the dimensionless KK–space, ranging from $1\times 10^{-3}$ to $3\times 10^{2}$. Each of the two hidden layers contains 512 neurons. The projection corresponding to this NN is strictly linear until an activation function is introduced. We adopt the Rectified Linear Unit (ReLU) as the activation function. Both the NN architecture and the activation function are selected based on our experimental results, which demonstrate their superior accuracy compared to other commonly adopted alternatives. It is worth noting that the NN does not take the phase transition temperature $T_{*}$ or the number of degrees of freedom $g_{*}$ as inputs, as the resulting gravitational wave spectrum today depends linearly on these parameters, as predicted by the enhanced SSM.
@@ 4.1 Estimating LISA noise and generating mock data
$\times\left(1+\left(\frac{f}{8\rm\,mHz}\right)^{4}\right)\left(\frac{1}{2\pi f}\right)^{4}\left(\frac{2\pi f}{c}\right)^{2}.$ (4.3)
They dominate at higher and lower frequencies, respectively, with $P\approx 15$, $A\approx 3$ as noise parameters. Based on this model, under the equilateral triangle assumptions of the three satellite and additional isotropic and stationary assumptions on $P_{\rm oms}(f,P)$ and $P_{\mathrm{a}cc}$  [^2] We refer to [38] for further details on noise estimations, the total noise in the measured energy fraction of GWs can be estimated as:
@@
The simulated signals $\Omega_{\rm GW}(f)$, used as test cases, are generated based on the enhanced SSM based on two representative sets of benchmark parameters. The benchmarks, $\mathrm{BP}_{1}$ and $\mathrm{BP}_{2}$, are listed in Table 2. They represent relatively strong phase transitions that occur during the electroweak epoch, resulting in a power-law-like signal and a comparatively flatter signal, respectively (see the blue curves in Fig. 3). The differences in their spectral shapes are mainly attributed to their distinct bubble wall velocities $v_{w}$. It is important to note that $\rm{BP}_{1}$ corresponds to the deflagration case with relatively strong phase transition strength. In this instance, due to the loss of kinetic energy, the GW content could be significantly suppressed compared to the predictions of the SSM. This discrepancy was first identified in [73] through hydrodynamical simulations and can be addressed by introducing a suppression factor into the SSM [74]. For now, this suppression has not been considered in the enhanced SSM code or the DeepSSM emulator. Therefore, caution should be made when employing this emulator, a suppression factor (such as the one proposed in [74]) may need to be manually introduced when exploring the parameter region with significant suppression. We plan to update both the enhanced SSM model and the emulator to accommodate this suppression effect in future work.
We generate the injected mock data using the GW spectra corresponding to these two benchmark points, calculated with the enhanced SSM code rather than the emulator. Any foreground signals, such as those introduced by galactic binaries, are approximately ignored in this work for simplicity. Techniques such as yearly modulation have proven highly effective in mitigating such foreground, particularly for power-law signals [75]. However, caution should be made when addressing peaked signals associated with phase transitions. As noted by [76], the broken power-law signal resulting from a FOPT can be approximately recovered for $\Omega_{\rm GW,peak}\geq 10^{-11}$ and $f_{\rm peak}\geq 2\times 10^{-3}\,\rm Hz$ from galactic binary foreground signals after yearly modulation, a threshold closely aligns with our benchmark points. Given the complex GW spectral shapes predicted by our model, when using our framework on signals contaminated by foreground noise, more careful studies are necessary to accurately extract phase transition parameters.
@@ 4.2 Bayesian inference
Additionally, studies such as [40] report a bimodal posterior for a benchmark point with $v_{w}=0.55$. However, a similar bimodal posterior is not observed in either of our benchmark points. This discrepancy may be attributed to differences in our benchmark phase transition parameters or to the difference models employed (an enhanced SSM-based emulator vs. an model combining SSM and an empirical template). Furthermore, multimodal posteriors are known to pose significant challenges for MCMC algorithms, as these methods often become trapped in isolated modes. Advanced techniques, such as multimodal nested sampling [81], may prove more effective in addressing such scenarios. We defer a comprehensive exploration of the parameter space and an investigation of potential multimodal degeneracies to future work.
The posteriors and contours presented in this section are based on the HMC sampler with a single chain. The entire sampling process typically requires 12,000 points to complete and takes less than 3 hours. During the MCMC sampling, the exceptional efficiency of the emulator allows the GW signal spectrum to be computed instantaneously, resulting in time costs being only attributed to likelihood computation and sampling processes. This large number of sampling point will make semi-analytical model such as the enhanced SSM impractical to be used in the MCMC, as each estimation typically needs 10 minuets to accomplished. Compared to Ref. [40], our framework demonstrates greater efficiency, likely attributable to the emulator’s high performance, the use of advanced samplers like HMC, and more cutting-edge hardware.
@@ 5 Conclusions
Note that the lifetime of sound waves $\tau_{\rm sw}$ is treated as a free parameter in our analysis. In principle, it should be closely related to other phase transition parameters (i.e., $\alpha$, $R_{*}$ and $v_{w}$). For example, $\tau_{\rm sw}$ can be approximated by the eddy turnover time, defined as the time before the fluid enters the non-linear regime $\tau_{\rm nl}\sim R_{*}/\sqrt{\Omega_{K}}$ [16, 28], where $\Omega_{K}$ is the kinetic energy fraction and can be calculated as shown in Refs. [67, 82, 83, 68, 84]. However, to maintain generality, we choose to consider $\tau_{\rm sw}$ as a free parameter when implementing the enhanced SSM, the emulator and the Bayesian inference. Thanks to the exceptional efficiency of our emulator, this approach may facilitate research on various choices of $\tau_{\rm sw}$. For instance, researchers are able to fix $\tau_{\rm sw}$ by incorporating their own estimates of $\tau_{\rm sw}$ when adopting our framework within their Bayesian pipelines.